How many permutations can be made from all the letters in the word "MISSISSIPPI", under the condition that "M" is not in the middle?


1
Expert's answer
2021-07-27T16:39:59-0400

Number of permutation in the word "MISSISSIPPI": "N_1=\frac{11!}{4!4!2!}=34650."

Number of permutations if "M" is in the middle: "N_2=\frac{10!}{4!4!2!}=3150."

Number of permutations if "M" is not in the middle: "N=N_1-N_2=31500."